# Real-Time Doppler and Ionospheric Dispersion Correction Techniques for Arbitrary Waveforms Utilizing GPU Compute

### v1

@@ II-B Numeric Doppler Dispersion Correction Via Interpolation
In this section, we present an overview analysis of various Doppler-dispersion correction methods, all of which seek to approximate the effect of (13) by numerically interpolating the transmitted signal. Table II lists the features of the Doppler dispersion correction methods considered in this work [^3] We determined the error drop off for windowed Whittaker–Shannon interpolation as well as frequency conversion numerically. The data presented in the table reflects the drop off for LFMs.. Of note, we highlight the Whittaker–Shannon interpolation (also known as sinc interpolation) and FFT P/Q resampling methods, both of which are exact for Nyquist-limited signals.
@@
TABLE II: Comparison of features of various Doppler dispersion correction/discrete resampling methods.
@@
Whittaker–Shannon Interpolation $\mathcal{O}\left(N^{2}\right)$ Exact Yes
Windowed Whittaker–Shannon Interpolation $\mathcal{O}\left(N\times N_{window}\right)$ $\mathcal{O}\left(N_{window}^{-2}\right)$ Yes
@@
In (16), $\Delta t$ is the sample period of the discrete signal. (16) is general and does not require upsampling to achieve exact performance. Unfortunately, the time complexity of Whittaker–Shannon interpolation scales as $\mathcal{O}\left(N^{2}\right)$. For more optimal interpolation rates, Whittaker–Shannon requires a window where only $N_{window}$ neighboring points are used in the summation. This was investigated in Ref. [21].
FFT P/Q resampling is mathematically identical to (16), but is performed in the frequency domain. In literature, FFT P/Q resampling can also be refered to as Whittaker–Shannon interpolation—we define these as separate algorithms for clarity. Via the Fourier transform, the sinc filter used in Whittaker–Shannon interpolation becomes a box filter, zeroing out or adding zeroed edge samples near $f=\pm\frac{f_{s}}{2}$. The size of this box filter is $\frac{N}{\alpha}$. After appending or removing the appropriate number of samples, an iFFT is performed using the same number of samples as the box filter, which interpolates the signal. This interpolation algorithm is common in image processing [22] as it is relatively simple to implement and can utilize optimized FFT libraries.
@@
A key drawback of FFT P/Q resampling vs. Whittaker–Shannon interpolation is that FFT P/Q resampling is only exact when $N-\frac{N}{\alpha}$ is an even integer. FFT P/Q resampling can be combined with frequency conversion for fine-tune adjustment between even-integer interpolant results. This is done by first performing FFT P/Q resampling to the nearest velocity that removes an even number of samples, and then performing frequency conversion to shift the remaining frequency difference. This presents a trade space for the optimal time dilation method that depends on the center frequency, bandwidth, signal length, and modulation.
III Analysis
In this section, we present analysis of the signal processing accuracy and rate of each algorithm. To compare the accuracy of each dispersion correction algorithm, we compared the arbitrary signal processing methods presented in section II with those presented in sections I-A and I-B. In order to compare algorithms directly, we performed an analysis of the signal processing accuracy via matched filtering of LFMs. To demonstrate the computational rates of the methods considered, all algorithms in this paper were implemented in CUDA 12.3 and performed on the NVIDIA H100 GPU with a $3.25~\text{GHz}$ CPU processor connected via PCIe 5.0. All calculations performed in both forms of analysis used a single-precision complex float, otherwise known as a “complex64”, for each I/Q sample. The signal processing accuracy and timing performance for the ionospheric dispersion correction are presented in section III-A. A similar analysis of the Doppler dispersion algorithms is presented in section III-B.
III-A Ionospheric Correction Performance Analysis
We tested (15) against the Pseudo–Chebyshev approximation method shown in (3). Each method was compared against truth data calculated via the cubic solution from (9) using trapezoidal integration to determine the phase. Results from one of our sample test cases can be seen in Fig. 2. As can be seen in this example, the FFT approximation method and the Pseudo–Chebyshev method resulted in losses on the order of $0.01~\text{dB}$ SNR. The difference between each method resulted in peaks separated by $0.001~\text{dB}$. This analysis indicated that (15) is nearly identical in accuracy to the Pseudo–Chebyshev polynomial approximation method for the parameters used to generate Fig. 2.
Figure 2: Comparison of the FFT-based ionospheric compensation method with both the undistorted waveform (labeled LFM), the Pseudo–Chebyshev polynomial approximation (labeled POLY), our FFT based method (labeled FFT), and our exact frequency from the cubic formula shown in Eq. 9 integrated via trapezoidal integration (labeled CUBIC). This test was run with $f_{0}=413~\text{MHz}$, $B=18~\text{MHz}$, a sample rate of $f_{s}=2.048~\text{GHz}$, $E=100e16~\text{electrons}/\text{m}^{2}$, and $T=100~\mu\text{s}$. In this particular example, $\text{CUBIC}\circledast\text{FFT}$ curve compares the FFT-based correction method with truth and lost $0.01~\text{dB}$ SNR. For this example, the Pseudo–Chebyshev polynomial expansion method compared similarly as close as $0.001~\text{dB}$, indicating that the FFT method is on par with the accuracy of the Pseudo–Chebyshev polynomial method.
The choice of GPU hardware allowed us to parallelize the FFT, iFFT, and the phase shifting of each sample in the receive window. Figure 3 shows the timing performance result of running (15) on a realistic waveform. Notably, Fig. 3b shows timing performance excluding the FFT and iFFT compute times. We separate these test cases to demonstrate the efficiency that can be gained from bundling the ionospheric correction with other digital signal processing methods that require computations in the frequency domain.
Figure 3: Histograms showing the results from $10^{7}$ time trials where we performed the algorithm shown in (15). This test was run with $f_{0}=413~\text{MHz}$, $B=18~\text{MHz}$, a sample rate of $f_{s}=2.048~\text{GHz}$, $E=100e16~\text{electrons}/\text{m}^{2}$, $T=100~\mu\text{s}$, and had $2^{19}$ total I/Q samples. Figure 3a shows the rates when performing ionospheric correction during pulse compression, which averaged about $46~\mu\text{s}$ to complete and correlates to a rate of approximately $11.3~\text{Gigasamples}/\text{s}$. Figure 3b excludes the FFT and iFFT times and averages approximately $2200~\text{ns}$, which correlates to about $240~\text{Gigasamples}/\text{s}$.
As can be seen in Fig. 3, on our hardware we were able to complete the ionospheric correction considered in about $46~\mu\text{s}$; nearly $44~\mu\text{s}$ of which are from performing the required FFTs. The signal processing time to compensate for ionospheric dispersion via (15) is significantly lower than the pulse width of the waveform itself, even at a relatively high sampling rate of $f_{s}=2.048~\text{GHz}$. This efficiency demonstrates the viability of (15) for use in real-time DSP.
III-B Performance Analysis of Doppler Dispersion Correction
An overview of the accuracy of each method from Table II is summarized in Fig. 4. For low velocities, the frequency conversion method has the lowest computational load and performs well. Even when at even sample removal, the FFT P/Q resampling with or without the tone has nearly equivalent accuracy to Whittaker–Shannon interpolation at the cost of variable compute times. At non-integer interpolants, the sinc/Whittaker–Shannon method performs the dilation with the least error.
Figure 4: Comparison of error as a function of target velocity for the various Doppler dispersion correction methods considered in Table II. The waveform used here had $f_{0}=411~\text{MHz}$, $B=18~\text{MHz}$, and $T=500~\mu\text{s}$. The waveform was originally sampled at a rate of $f_{s0}=2~\text{GHz}$, and was downconverted to $f_{s}=200~\text{MHz}$. Notice that the the error for FFT P/Q resampling is harmonic, with no error at velocities that require an even integer number of samples be padded and maximal error for velocities that require an odd integer number of samples be added. Note that for the sinc interpolation loss shown here, a window size of $N_{window}=25~\text{samples}$ was used as well as a lookup table (LUT) that was sampled such that for $\operatorname{sinc}(x)$, $\Delta x=1/100$ ($100~\text{values}/\text{sample}$). Refer to Appendix C for details on the implementation of LUTs.
For FFT P/Q resampling, the signal can be resampled during an FFT by padding the signal with an even number of samples. When the number of interpolants, $N_{I}$, is a power of two, FFT P/Q resampling completes in $\mathcal{O}\left(N_{I}\log{N_{I}}\right)$ time, but the time complexity is much more sophisticated and almost always less optimal in other cases. The variable time complexity is demonstrated in Fig. 5 where three separate clusters of Doppler dispersion correction rates can be observed, each cluster correlated to a range of target velocities. This means that FFT P/Q resampling has a pseudo-velocity-dependent time complexity, which can result in unpredictable signal processing rates. The test case used in Fig. 5 is considered relatively stressful in terms of typical real-time DSP, and the required compute could be considered viable for some systems. While the simplicity of implementing FFT P/Q resampling is desirable, the variable time complexity is likely only be valid for systems with low sample rates or range extents.
@@ III-B Performance Analysis of Doppler Dispersion Correction
Figure 5: Doppler dispersion correction rates for $10^{6}$ trials of FFT P/Q Resampling with a waveform such that $f_{0}=420~\text{MHz}$, $B=18~\text{MHz}$, $T=500~\mu\text{s}$, and $N=2^{19}$. To demonstrate the variable time complexity, we also varied the target velocity by using a uniform distribution between $v=0\text{ to }5~\text{km}/\text{s}$. Notice three separate clusters of completion times, each related to a different range of velocities. The left-most cluster is a sharp peak with times $\leq 10~\mu\text{s}$ and corresponds to about $v<1500~\text{km}/\text{s}$ where no samples were added (no work was done). The right-most peak was for velocities such that about $1500<v<4500~\text{km}/\text{s}$, where only 2 samples were removed. The final middle cluster was the remaining trials during which 4 samples were removed. The average completion rates for the left, center, and right cluster are about $6.5~\text{Terasamples}/\text{s}$, $670~\text{Megasamples}/\text{s}$, and $260~\text{Megasamples}/\text{s}$ respectively.
@@
Considerations for efficient implementation of Whittaker–Shannon interpolation are presented in Appendix C. Our optimal implementation of the algorithm is presented in Alg. 2 and resulted in interpolation rates shown in Fig. 6 of about $11.7~\text{Gigasamples}/\text{s}$. For a radar transmitting at 10% duty factor, this would mean a rate of $117~\text{Gigasamples}/\text{s}$ could be sampled on receive.
@@
Figure 6: timing performance measurements of Whittaker–Shannon interpolation times on an NVIDIA H100 GPU for $10^{7}$ time trials. The input waveform contained $N=2^{19}~\text{samples}$ total with a window size of $N_{window}=25~\text{samples}$. Interpolation was achieved at an average of $45~\mu s$, or a rate of about $11.7~\text{Gigasamples}/s$.
@@ IV Conclusion
In this paper, we investigated potential algorithms for ionospheric and Doppler dispersion correction of arbitrary waveforms. The goal of this analysis was to identify algorithms which are viable for real-time DSP on GPU hardware, allowing for simpler system designs. We also desired flexible algorithms by implementing numerical approximation methods that could be applied directly to discrete signals with no requisite knowledge of the transmitted waveform.
For ionospheric dispersion correction, we proposed an FFT-based method which was presented in (15). This algorithm is computed directly from the ionospheric delay shown in (1). The primary benefits of this method are that it requires no previous information of the input waveform to compute, is valid for both predistortion and post-distortion, is valid for arbitrary waveforms, and is straightforward to implement. The major drawback of this algorithm is the relatively high computational cost, which can only be mitigated with more powerful parallel compute hardware. This algorithm also has an SNR loss of $\mathcal{O}(t_{total})$, which means that the accuracy is reduced for smaller range window sizes. The signal processing accuracy of (15) was shown in Fig. 2 on LFMs. This demonstrated that the algorithm was approximately as accurate as the Pseudo–Chebyshev polynomial expansion method, which indicates a suitable accuracy for radar DSP. In Figs. 3a and 3b we provided timing performance of an implementation of (15) on an NVIDIA H100 GPU. These timings demonstrate that this method can be viable for application in real-time DSP given appropriate selection of hardware and parallelization.
We also presented a comparison of various Doppler dispersion correction methods via time dilation. We highlighted the methods of FFT P/Q resampling and windowed Whittaker–Shannon interpolation, given in (16), as potentially viable candidates for Doppler dispersion correction in SDRs. Figure 4 presents the accuracy of various Doppler compensation algorithms considered. The data processing rates of FFT P/Q resampling and sinc interpolation on GPU hardware are shown in Figures 5 and 6 respectively. Both algorithms were demonstrated to have sufficient accuracy to be considered for use in radar DSP. The signal processing rates of windowed Whittaker–Shannon interpolation were much more consistent and rapid than those of FFT P/Q resampling. While windowed Whittaker–Shannon interpolation was shown to be valid for radar DSP, the variable processing rates of FFT P/Q resampling mean that it is likely only viable for systems with low sampling rates or range window sizes until modern GPU hardware improves. The major benefit of FFT P/Q resampling is its simple implementation and optimization, which results in relatively little code. On the other hand, optimization of efficient windowed Whittaker–Shannon interpolation is more challenging. To reduce this challenge, we provide considerations for optimizing windowed Whittaker–Shannon interpolation in Appendix C. Despite these considerations, the ability to implement windowed Whittaker–Shannon interpolation on a GPU simplifies system design time relative to radar-specific hardware. We conclude that windowed Whittaker–Shannon interpolation is the most-viable candidate for real-time radar DSP considered in this work and would be sufficient for many modern real-time DSP applications.
@@ Appendix A Ionospheric Effects on Radio Signals
$n_{G}=1+\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}.$ (17)
Here, $N_{e}$ is the ionospheric electron density and $\epsilon_{0}$ is the permittivity of free space. Thus, the group velocity of an ionospheric radio wave, $V_{G}$, is given as (18):
$\begin{split}V_{G}=\frac{c}{n_{G}}&=c\left(1+\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}\right)^{-1}\\ &=c\left(1-\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}\right)+\mathcal{O}\left(f^{-4}\right).\end{split}$ (18)
@@
$R_{e}=\int_{\Delta t}V_{G}dt=c\Delta t-\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}\int_{\Delta t}cN_{e}dt.$ (19)
@@
$R_{e}=c\Delta t-\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}\int_{s}N_{e}ds.$ (20)
@@
$\begin{split}\tau(f)&=\frac{R_{0}-R_{e}}{c}=\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}\int_{s}N_{e}ds\\ &=\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}E=\frac{K_{2}}{cf^{2}}.\end{split}$ (21)
Appendix B Correction to Pseudo–Chebyshev Polynomial Expansion Method
The original Pseudo–Chebyshev method presented in Ref. [1] slightly differed from the coefficients we presented in (7) and (8). The original forms of $T_{1}$ and $T_{2}$ were given as:

### v2

@@ III-A Doppler Dispersion Time Dilation Methods
In this section, we present an overview analysis of various Doppler-dispersion correction methods, all of which seek to approximate the effect of (2). Most methods will approximate (2) via a numeric interpolation of the transmitted signal. Table I lists the Doppler dispersion correction methods considered here along with some basic features of each method [^2] We determined the error drop off for windowed Whittaker-Shannon interpolation as well as frequency conversion numerically. The data presented in the table is reflective of that drop off for LFMs..
Whittaker-Shannon interpolation (also known as sinc interpolation) is an exact interpolation method for Nyquist limited signals, and is given by (17):
@@
TABLE I: Comparison of features of various Doppler dispersion correction/discrete resampling methods.
@@
Whittaker-Shannon Interpolation $\mathcal{O}\left(N^{2}\right)$ Exact Yes
Windowed Whittaker-Shannon Interpolation $\mathcal{O}\left(N\times N_{window}\right)$ $\mathcal{O}\left(N_{window}^{-2}\right)$ Yes
@@
Here, $\Delta t$ is the sample period of the discrete signal. (17) is general and does not require upsampling to achieve exact performance. Unfortunately, the time complexity of Whittaker-Shannon interpolation scales as $\mathcal{O}\left(N^{2}\right)$. For more optimal interpolation rates, Whittaker-Shannon requires a window where only $N_{window}$ neighboring points are used in the summation. This was investigated in Ref. [15].
@@
For FFT P/Q resampling, the signal can be resampled during an FFT by padding the signal with an even number of samples. When the number of interpolants, $N_{I}$, is related to the number of input samples by a factor of a power of 2, FFT P/Q resampling completes in $\mathcal{O}\left(N_{I}\log{N_{I}}\right)$ time, but the time complexity is much more sophisticated and almost always less optimal in other cases. The variable time complexity is demonstrated in Fig. 4 where three separate clusters of Doppler dispersion correction rates can be observed, each cluster relating to a range of target velocities. When the desired number of interpolants is not an even integer, FFT P/Q resampling is not exact. FFT P/Q Resampling can be combined with frequency conversion for fine-tune adjustment between even-integer interpolant results. This is done by first performing FFT P/Q Resampling to the nearest velocity that removes an even number of samples, and then performing frequency conversion to shift the remaining frequency difference. This presents a trade space for the optimal time dilation method that depends on the center frequency, bandwidth, signal length, and modulation.
@@
Figure 4: Doppler dispersion correction rates for $10^{6}$ trials of FFT P/Q Resampling with a waveform such that $f_{0}=420~\text{MHz}$, $B=18~\text{MHz}$, and $T=500~\mu\text{s}$. To demonstrate the variable time complexity, we also varied the target velocity by using a uniform distribution between $v=0\text{ to }5~\text{km}/\text{s}$. Notice three separate clusters of completion times, each related to a different range of velocities. The left-most cluster is a sharp peak with times $\leq 10~\mu\text{s}$ and corresponds to about $v<1500~\text{km}/\text{s}$ where no samples were added (no work was done). The right-most peak was for velocities such that about $1500<v<4500~\text{km}/\text{s}$, where only 2 samples were removed. The final middle cluster was the remaining trials during which 4 samples were removed. The average completion rates for the left, center, and right cluster are about $6.5~\text{Terasamples}/\text{s}$, $670~\text{Megasamples}/\text{s}$, and $260~\text{Megasamples}/\text{s}$ respectively.
An overview of the accuracy of each method is summarized in Fig. 5. For small dilations, the frequency conversion method has the lowest computational load and performs well. Even when at even sample removal, the FFT P/Q resampling with or without the tone has nearly equivalent accuracy to Whittaker-Shannon interpolation at the cost of variable compute times. At non-integer interpolants, the sinc/Whittaker-Shannon method performs the dilation with the least error.
Figure 5: Comparison of error as a function of target velocity for the various Doppler dispersion correction methods considered in Table I. The waveform used here had $f_{0}=411~\text{MHz}$, $B=18~\text{MHz}$, and $T=500~\mu\text{s}$. The waveform was originally sampled at a rate of $f_{s0}=2~\text{GHz}$, and was downconverted to $f_{s}=200~\text{MHz}$. Notice that the the error for FFT P/Q resampling is harmonic, with no error at velocities that require an even integer number of samples be padded and maximal error for velocities that require an odd integer number of samples be added. Note that for the sinc interpolation loss shown here, a window size of $N_{window}=25~\text{samples}$ was used as well as a lookup table (LUT) that was sampled such that for $\operatorname{sinc}(x)$, $\Delta x=1/100$ ($100~\text{values}/\text{sample}$). Refer to Appendix C for details on the implementation of LUTs.
III-B Implementation of Whittaker-Shannon Interpolation
Whittaker-Shannon interpolation requires a new FIR filter be calculated at every time-step. For each sample, a local filter is calculated using the interpolation shift at the point. This filter will always be whole sample steps plus a fine adjustment when the sampling is between points. Using analytic resampling as a control case for comparison of the windowed Whittaker-Shannon interpolation method, a comparison of the accuracy of sinc interpolation to an exact solution can be seen in Fig. 6.
Figure 6: Comparison of windowed Whittaker-Shannon interpolation with the exact analytic interpolation method. This image shows the SNR loss from using sinc interpolation on an LFM waveform with a pulse width $T=0.1~\text{ms}$, sample rate of $f_{s}=2.048~\text{GHz}$, a center frequency of $f_{c}=422~\text{MHz}$, bandwidth of $B=18~\text{MHz}$, and a target velocity of $v=1~\text{km}/\text{s}$. Here the DIR curve represents the analytic solution for Doppler dispersion correction via direct resampling and SINC represents the result of Doppler compensating via windowed Whittaker-Shannon interpolation with a window size of $N_{window}=41~\text{samples}$. The resultant difference in SNR is about a $10^{-4}~\text{dB}$ SNR loss.
@@ III-B Implementation of Whittaker-Shannon Interpolation
Considerations for efficient implementation of sinc interpolation are presented in Appendix C. Our implementation of the code resulted in interpolation rates of about $11.7~\text{Gigasamples}/\text{s}$, and a set of benchmarking data can be seen in Fig. 7. This interpolation only has to be applied for the matched filter. For a radar transmitting at 10% duty factor, this would mean a rate of $117~\text{Gigasamples}/\text{s}$ could be sampled on receive.
@@
Figure 7: Benchmarks of Whittaker-Shannon interpolation times on an NVIDIA H100 GPU for $10^{7}$ time trials. The input waveform contained $N=2^{19}~\text{samples}$ total with a window size of $N_{window}=25~\text{samples}$. Interpolation was achieved at an average of $45~\mu s$, or a rate of about $11.7~\text{Gigasamples}/s$.
@@ IV Conclusion
In this paper, we propose an ionospheric correction algorithm for digital processing with SDR, presented in (16). Our algorithm is computed directly from the ionospheric delay shown in (3). The primary benefits of our method are that it requires no previous information of the input waveform to compute, is valid for both pre-distortion and post-distortion, is valid for arbitrary waveforms, and is straight-forward to implement. Shown in Fig. 2, we compare our proposed method to a common approximate solution for application on LFMs as well as with an exact analytic method we derived and is shown in (11). In doing so, we demonstrate an SNR loss within the error tolerance of the known approximation method, which indicates satisfactory accuracy. In Figs. 3(a) and 3(b), we provide benchmarks of an implementation of our algorithm on an NVIDIA H100 GPU. These benchmarks demonstrate that our method can be viable for application in modern digital radar given appropriate selection of hardware and parallelization.
We also presented a comparison of various Doppler dispersion correction methods via time-dilation interpolation. We highlighted the method of windowed Whittaker-Shannon interpolation as a viable candidate for Doppler dispersion correction in SDRs. Figures 6 and 7 present examples of the accuracy and data processing rates respectively of sinc interpolation in software. We conclude that windowed Whittaker-Shannon interpolation is a viable candidate for applications in real-time DSP.
@@ Appendix A Ionospheric Effects on Radio Signals
$n_{G}=1+\frac{e^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}.$ (18)
Here $N_{e}$ is the ionospheric electron density, $\epsilon_{0}$ is permittivity of free space, and $e$ and $m_{e}$ are the charge and mass of an electron, respectively. Thus, the group velocity of an ionospheric radio wave, $V_{G}$, is given as (19):
$\begin{split}V_{G}=\frac{c}{n_{G}}&=c\left(1+\frac{e^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}\right)^{-1}\\ &=c\left(1-\frac{e^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}\right)+\mathcal{O}\left(f^{-4}\right).\end{split}$ (19)
@@
$R_{e}=\int_{\Delta t}V_{G}dt=c\Delta t-\frac{e^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}\int_{\Delta t}cN_{e}dt.$ (20)
@@
$R_{e}=c\Delta t-\frac{e^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}\int_{s}N_{e}ds.$ (21)
@@
$\begin{split}\tau(f)&=\frac{R_{0}-R_{e}}{c}=\frac{e^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}\int_{s}N_{e}ds\\ &=\frac{e^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}E=\frac{K_{2}}{cf^{2}}.\end{split}$ (22)
Appendix B Correction to Pseudo-Chebyshev Polynomial Expansion Method
The original Pseudo-Chebyshev method presented in Ref. [1] slightly differed from the coefficients we presented in (9) and (10). The original forms of $T_{1}$ and $T_{2}$ were given as: